find an equation for the line with the given properties. Express the equation in slope-intercept form
slope=6; containing the point (3,-8)

Answer by mananth(16949)

slope m = (y2-y1)/(x2-x1)
( 1 - -2 )/( -2 - 1 )
( 3 / -3 )
m= 6

Plug value of the slope and point ( 3,-8 ) in
Y = m x + b
-8.00 = 18 + b
b= -8 - 18
b= -26
So the equation will be
Y =6x-26
m.ananth@hotmail.ca